A 0 20-4m wide river has a uniom flow speed of 4.054 m/s toward the east A boat with a speed of 80 m/s relative to the water leaves the south bank and heads in such a way that it crosses to a point directly north of its depaiture point How long does it take the boat to cross the river? 295 23s 0 25 s 0 27s 0 17s
Answer #1

To counter the river's drift velocity, the boat will have a component of velocity of the magnitude of 4.054m/s to the west. The velocity component from South to North can be calculated using vector addition formula

v_{SN} = \sqrt{8^2 - 4.054^2} = 6.89m/s

Therefore the time is given by

t = \frac{d}{v_{SN}} = \frac{200}{6.89} = 29m/s

Please upvote if the answer was useful. Comment for doubts and clarifications.